# Harmony in Divergence: Towards Fast, Accurate, and Memory-efficient Zeroth-order LLM Fine-tuning

### v1

@@ 3.3 Projection Learning by Zeroth-order Optimization
^$\nabla\widehat{\mathcal{L}}(\bm{\gamma}_{t};\bm{\theta}_{t})=\left[\frac{\widehat{\mathcal{L}}\left(\bm{\gamma}_{t}+\epsilon\bm{u};\bm{\theta}_{t}\right)-\widehat{\mathcal{L}}\left(\bm{\gamma}_{t}-\epsilon\bm{u};\bm{\theta}_{t}\right)}{2\epsilon}\bm{u}\right]$ (6)
@@
where $\bm{u}\in\mathbb{R}^{L}$ is a random vector from $\mathcal{N}(0,\mathbf{I})$, $\widehat{\mathcal{L}}$ is the objective defined in Eq. (4).
However, naively applying vanilla ZO optimization for the sub-optimization (projection learning) results in unsatisfactory enhancement. More critically, it can lead to sub-optimization failure and undermine the main fine-tuning process (see Appendix C.2 for results). Two primary issues contribute to the failure. First, the values of projections are not only related to $\bm{\gamma}_{t}$ but also the distance gap $\|\Delta\bm{\theta}_{t}\|$. Ignoring the distance gap when searching for projections causes uninformative optimization and yields sub-optimal solutions. Second, because the projection is derived through noisy ZO optimization over only a few iterations, there is a risk that the projection magnitude becomes inappropriately small or large. A small projection drives the fine-tuned model too close to the pre-trained model, nullifying many previous updates, while a large projection applies overly aggressive weight updates, destabilizing the training process.
@@
Re-initialization. To introduce the distance gap $\|\Delta\bm{\theta}_{t}\|$ into the projection learning process, the initial value $\bm{\gamma}_{t,0}$ is reset to $\|\Delta\bm{\theta}_{t}\|$ each time the projection is optimized. This means that, initially, the projection magnitude $\frac{\bm{\gamma}_{t}}{\|\Delta\bm{\theta}_{t}\|}=1$. If projection updates are not performed, DiZO reverts to standard ZO optimization.
@@
With the above two strategies, we enhance the learning process of projection, more analysis can be found in Appendix C.2. We also provide a Pytorch-style implementation, please refer to Appendix F for details.
4 Discussion and Overhead Analysis
@@
We have some discussion on our method and analyze the computational overhead here and elaborate further later.
Would adjusting the learning rate be equally effective? As discussed in Section 2.2, our main objective is to provide ZO optimization with diverse-magnitude updates. A seemingly straightforward alternative is to assign different learning rates to each layer. However, in practice, this approach yields results that are similar to or even worse than vanilla ZO in terms of accuracy and GPU hours. We attribute this to the noisy gradient estimation of one single ZO step, which is likely to yield imprecise update directions. Therefore, using unrefined layer-wise learning rates can intensify this noise and further destabilize the optimization process. In contrast, DiZO enables the awareness of the pre-trained model during fine-tuning (see Eq. 5), robustifies the training process (Dong et al., 2021; Oh et al., 2023; Zhai et al., 2023; Wang et al., 2024), and mitigates the noise introduced by ZO’s random perturbations. More results and analysis are shown in Appendix C.3.
@@
Memory utilization. Our method requires additional memory as it involves storing the pre-trained model and calculating the weight distance gap with the fine-tuned model, which can become costly when scaling to large LLMs. However, in DiZO, we find that projecting only the weight updates of the Query and Value layers in the attention module, instead of updating all layers, not only reduces memory requirements but also delivers better performance. As a result, we only need to store the weights of these two types of layers from the pre-trained model, accounting for approximately 16.7% of the parameters in OPT-2.7B, which is a manageable overhead. Similarly, LoRA (Hu et al., 2021) also focuses on weight decomposition for Query and Value layers, which echoes our observation. Further analysis and results on projection layer selection are provided in Appendix C.1.
@@
Computational overhead. Our method introduces extra computational cost, as the projection is learned alongside the main optimization (fine-tuning). However, we observe that performing projection learning intermittently, only once every few training iterations, does not compromise performance and significantly reduces the added overhead. This strategy reduces the computational burden while maintaining efficiency, allowing DiZO to achieve throughput comparable to vanilla ZO fine-tuning. Additionally, the reduced frequency of projection updates ensures that DiZO remains scalable for larger models and datasets. Please refer to Section 5.4 and Appendix D.1 for more details on computational overhead.
@@ 5 Experiments
Table 3: Experiment results on RoBERTa-large (350M) on six classification datasets. Results of the baseline methods MeZO and MeZO LoRA are taken from Malladi et al. (2023). All reported numbers are averaged accuracy with standard deviation shown. Better results between MeZO and DiZO are highlighted in bold.
Dataset Task Type SST-2 SST-5 SNLI MNLI RTE TREC
——-sentiment——- ———-language inference———- –topic–
Zero-shot 79.0 35.5 50.2 48.8 51.4 32.0
Gradient-free methods: $k=16$
MeZO 90.5 (1.2) 45.5 (2.0) 66.0 (2.7) 56.5 (2.5) 59.4 (5.3) 76.9 (2.7)
DiZO 92.2 (0.9) 47.1 (1.3) 71.0 (3.1) 60.1 (3.5) 67.9 (4.7) 77.4 (2.4)
Gradient-based methods: $k=16$
FT 91.9 (1.8) 47.5 (1.9) 77.5 (2.6) 70.0 (2.3) 66.4 (7.2) 85.0 (2.5)
Gradient-free methods: $k=512$
MeZO 93.3 (0.7) 53.2 (1.4) 83.0 (1.0) 78.3 (0.5) 78.6 (2.0) 94.3 (1.3)
DiZO 94.6 (0.1) 53.6 (1.7) 84.5 (0.6) 79.8 (0.9) 80.3 (1.8) 93.8 (1.5)
Gradient-based methods: $k=512$
FT 93.9 (0.7) 55.9 (0.9) 88.7 (0.8) 84.4 (0.8) 82.7 (1.4) 97.3 (0.2)
Figure 2: Trajectory of training loss curves when using MeZO and DiZO to fine-tune Roberta-large on SST-2, MNLI, and RTE.
@@ 5.1 Experimental Settings
Models and datasets. We evaluate DiZO with various models, including medium-sized masked models (Liu et al., 2019) (RoBERTa-large) and large-sized autoregressive models (Zhang et al., 2022; Touvron et al., 2023) with different size, including OPT-2.7B, OPT-6.7B, Llama3-3B, and Llama3-8B. The total parameter size is ranging from 355M to 8B. Both classification and generation tasks are included. More details on datasets are shown in Appendix B.1.
Baseline. We mainly compare with two ZO works, memory-efficient ZO optimization (MeZO) (Malladi et al., 2023) and Hessian-informed ZO optimization (HiZOO) (Zhao et al., 2024). MeZO is a fundamental and representative work in ZO LLM fine-tuning but suffers from slow convergence speed. HiZOO [^1] We implement HiZOO ourselves, please refer to Appendix B.2 for details. is a recently proposed ZO acceleration for LLM fine-tuning, which leverages the estimated second-order information to speed up. In addition, we also incorporate the parameter-efficient fine-tuning (PEFT) technique LoRA (Hu et al., 2021), applying it on top of FO fine-tuning, MeZO, and HiZOO.
Evaluation. For training and evaluation, we follow previous works (Gao et al., 2020; Malladi et al., 2023). We study few-shot and many-shot settings on RoBERTa-large, randomly sampling $k$ samples per class for training and validation, and 1000 samples for testing. For RoBERTa models, we evaluate $k=16$ and $k=512$. For OPT and LLaMA, we sample 1000, 500, and 1000 samples for training, validation, and testing. All experiments are conducted on NVIDIA A100 and A6000 GPUs.
@@ 5.2 Medium-sized masked language models
We conduct experiments on RoBERTa-large across three types of datasets and compare DiZO with two ZO baselines. We also explore PEFT by integrating LoRA. Table 3 presents the results, while Figure 2 shows the trajectory of training loss curves, indicating the convergence speed of DiZO and MeZO. Our key findings are as follows:
DiZO greatly increases the convergence speed over MeZO. By using divergence-driven layer adaptation, the loss curve of DiZO decreases much faster, cutting the required iterations by over 50% on SST-2, MNLI, and RTE. In addition, DiZO improves accuracy by 1.7%, 3.6%, and 8.5% on these three datasets, respectively.
DiZO outperforms MeZO and achieves results on par with full fine-tuning. From Table 3, DiZO consistently surpasses MeZO on all six datasets. Notably, on SST-2 and RTE datasets, DiZO even shows better performance than FO full-parameter fine-tuning, increasing by 0.3% and 1.5%, respectively.
@@ 5.3 Large autoregressive language models
Table 4: Experiments results of fine-tuning OPT-2.7B on seven classification datasets and two text generation datasets (with 1000 training samples). Better results between MeZO, HiZOO, and DiZO are highlighted in bold.
@@
Table 5: Experiment results on OPT-6.7B for four classification datasets and one text generation dataset (with 1000 training samples). Better results are highlighted in bold.
Dataset Task Type SST-2 RTE CB WSC SQuAD
@@
Figure 3: Experiment result on Llama3-3B and Llama3-8B for four classification datasets and one text generation dataset. More results and detailed numbers are shown in Appendix D.2.
To assess the broader applicability of DiZO, we run experiments on the OPT and Llama series autoregressive LLMs covering both classification and generation tasks. The overall results are summarized in Table 4, Table 3, and Figure 3 for OPT-2.7B, OPT-6.7B, and Llama series, respectively. We also compare the convergence speeds of DiZO and MeZO on OPT-2.7B across multiple datasets in Figure 4. Below, we highlight the key observations from these experiments.
DiZO dramatically reduces the training GPU hours compared with the representative baseline MeZO. By incorporating divergence-driven layer adaptation, DiZO quickly establishes meaningful divergence across layers, whereas MeZO requires many more iterations to achieve the desired layer-wise divergence. As shown in Table 4, DiZO converges with far fewer iterations across nine datasets, resulting in up to a 48% reduction in training GPU hours. Moreover, unlike HiZOO, which reduces the number of iterations needed but slows the throughput of MeZO by more than 1.5× due to Hessian estimation, DiZO keeps its throughput nearly on par with MeZO. This efficiency is achieved because the additional projection learning procedure needs only two forward passes and is performed intermittently.
DiZO outperforms baselines in both standard and parameter-efficient settings. From Table 4, DiZO surpasses MeZO and HiZOO with or without the LoRA, achieving results comparable to FO methods. Across seven classification datasets, DiZO ranks first on five, and it also leads in both text generation tasks. Table 3 shows that these advantages persist even when scaling up to OPT-6.7B. Moreover, as illustrated in Figure 3, the fine-tuning process of Llama-series model also benefits from layer-wise adaptive updates.
Figure 4: Comparison of convergence iterations, forward pass, and training GPU hours between MeZO and DiZO across multiple datasets. Results are presented as proportions, with the percentage of saved GPU hours highlighted for each dataset.
@@ 5.4 Memory and Speed Analysis
Table 6: Memory utilization and speed test on OPT-2.7B on RTE dataset (180 tokens per example on average). : partial gradient-free; : gradient-free; : gradient-based. DiZO: learning projection with Adam. For a fair comparison, the speed and memory are measured on the same machine with the same setting using the same batch size. Please refer to Appendix D.1 for results on more datasets.
@@
In this section, we examine the memory utilization and convergence speed of DiZO in comparison with both ZO baselines and FO fine-tuning approaches (with and without LoRA). Table 6 presents the results of fine-tuning OPT-2.7B on the RTE dataset, more results are shown in Appendix D.1.
From the memory perspective, DiZO maintains the advantage of avoiding backpropagation, getting rid of the storage of memory-intensive data, and reducing memory usage by about 90% compared to FO fine-tuning. As explained in Section 4, the additional memory requirement of DiZO stems from storing a portion of the pre-trained weights, including the weight of the Query and Value, amounting to only 16.7% of the total parameters. In contrast, HiZOO needs to store Hessian information for all layers, with memory usage proportional to the size of the parameters.
From the perspective of convergence speed, DiZO greatly reduces the required iterations while maintaining throughput similar to MeZO, resulting in significantly fewer training GPU hours. In contrast, HiZOO does not achieve comparable iteration savings and lowers the throughput of MeZO by about 1.5× because it requires Hessian information estimation. As a result, it only shows a modest acceleration in training GPU hours, in some settings, such as HiZOO with LoRA on RTE, it even consumes more training GPU hours than MeZO with LoRA.
@@
A notable byproduct of our method is using a FO approach (e.g., with the Adam optimizer) to learn the projections. While this version has memory consumption comparable to LoRA and requires additional training GPU hours, it offers distinct advantages. Since DiZO does not update projections at every iteration, FO-based DiZO exhibits significantly lower average memory usage than FO-based LoRA, with an average memory overhead close to that of the ZO-based DiZO. Although average memory usage may seem less critical in single-process, single-GPU setup, many real-world on-device training scenarios involve multi-process environments (Li et al., 2024; Ye et al., 2024). In such cases, the FO-based DiZO can stagger memory usage phases across processes, enabling parallel operations that purely FO methods cannot achieve. Furthermore, compared with ZO-based DiZO, the FO version reduces extra training GPU hours and delivers better performance. These qualities make it particularly appealing for specific on-device training cases.

### v3

@@ 3.3 Projection Learning by Zeroth-order Optimization
$\nabla\widehat{\mathcal{L}}(\bm{\gamma}_{t};\bm{\theta}_{t})=\left[\frac{\widehat{\mathcal{L}}\left(\bm{\gamma}_{t}+\epsilon\bm{u};\bm{\theta}_{t}\right)-\widehat{\mathcal{L}}\left(\bm{\gamma}_{t}-\epsilon\bm{u};\bm{\theta}_{t}\right)}{2\epsilon}\bm{u}\right]$ (6)
@@
where $\bm{u}\in\mathbb{R}^{L}$ is a random vector from $\mathcal{N}(0,\mathbf{I})$ and $\widehat{\mathcal{L}}=\mathcal{L}(\bm{\theta_{0}}+\frac{\bm{\gamma_{t}}}{\|\Delta\bm{\theta_{t}}\|}\Delta\bm{\theta_{t}};\mathcal{B}_{t})$.
However, directly applying vanilla ZO optimization to the sub-task of projection learning yields limited improvement and can even cause failure, undermining the main fine-tuning process (see Appendix C.2). This failure stems from two key issues. First, projection values depend not only on $\bm{\gamma}_{t}$ but also on the distance gap $\|\Delta\bm{\theta}_{t}\|$. Ignoring this gap leads to uninformative updates and suboptimal solutions. Second, due to noisy ZO updates over a few iterations, projection magnitudes can become too small or too large. A small projection pulls the model too close to the pre-trained state, erasing progress, while a large one applies overly aggressive updates that destabilize training.
@@
Re-initialization. To introduce the distance gap $\|\Delta\bm{\theta}_{t}\|$ into the projection learning process, the initial value $\bm{\gamma}_{t,0}$ is reset to $\|\Delta\bm{\theta}_{t}\|$ each time the projection is optimized. This means that, initially, the projection magnitude $\frac{\bm{\gamma}_{t}}{\|\Delta\bm{\theta}_{t}\|}=1$. If not perform projection updates, DiZO reverts to standard ZO.
@@
4 Overhead Analysis
@@
We simply analyze the computational overhead of our method here and will elaborate further later.
@@
Memory overhead. Our method requires additional memory as it involves storing the pre-trained model and calculating the weight distance gap with the fine-tuned model, which can become costly when scaling to large LLMs. However, in DiZO, we find that projecting only the weight updates of the Query and Value layers in the attention module, instead of updating all layers, not only reduces memory requirements but also delivers better performance. As a result, we only need to store the weights of these two types of layers from the pre-trained model, accounting for approximately 16.7% of the parameters in OPT-2.7B, which is a manageable overhead. Similarly, LoRA hu2021lora also focuses on weight decomposition for Query and Value layers, which echoes our observation.
@@
Computational overhead. Our method introduces extra computational cost, as the projection is learned alongside the main optimization (fine-tuning). However, we observe that performing projection learning intermittently, only once every few training iterations, does not compromise performance and significantly reduces the added overhead. This strategy reduces the computational burden while maintaining efficiency, allowing DiZO to achieve throughput comparable to vanilla ZO fine-tuning. Additionally, the reduced frequency of projection updates ensures that DiZO remains scalable for larger models and datasets.
5 Convergence Analysis
In this section, we give a nonconvex convergence guarantee in terms of the expected gradient norm. The bound improves over basic ZO-SGD by replacing the factor $d$ (full dimension) with an effective dimension on the order of $D$ where we denote by $D=\max_{1\,\leq\,\ell\,\leq\,L}\,d^{(\ell)}$.
We assume the following, which are standard in ZO analyses:
Assumption 5.1.
$\mathcal{L}$ is $L_{f}$-smooth, i.e. there exists $L_{f}>0$ such that for all $\bm{\theta},\bm{\theta}^{\prime}$,
$\|\nabla\mathcal{L}(\bm{\theta})\;-\;\nabla\mathcal{L}(\bm{\theta}^{\prime})\|\;\leq\;L_{f}\,\|\bm{\theta}-\bm{\theta}^{\prime}\|.$
Assumption 5.2.
Due to DiZO’s layerwise projection step, each $\bm{\theta}^{(\ell)}$ remains in a ball (or line segment) around $\bm{\theta}_{0}^{(\ell)}$ of radius $R^{(\ell)}$. In particular, define
$\mathcal{S}\;=\;\bigl\{\bm{\theta}\,\mid\,\|\bm{\theta}^{(\ell)}-\bm{\theta}_{0}^{(\ell)}\|\,\leq\,R^{(\ell)}\;\;\forall\,\ell\bigr\}.$
The algorithm ensures $\bm{\theta}_{t}\in\mathcal{S}$ for all iterations $t$.
Theorem 5.3.
Under Assumptions 5.15.2, suppose DiZO runs for $T$ iterations with step size $\eta=c/\sqrt{T}$ for a sufficiently small constant $c>0$. Then there exist constants such that
$\min_{0\,\leq\,t\,<\,T}\mathbb{E}\bigl[\|\nabla\mathcal{L}(\bm{\theta}_{t})\|^{2}\bigr]\;\;=\;\;O\!\Bigl(\frac{\sqrt{D}}{\sqrt{T}}\Bigr).$
Note that $D$ can often be much smaller than the total parameter count $\sum_{\ell=1}^{L}d^{(\ell)}$. This drastically improves the variance bounds in the zeroth-order gradient estimation, leading to a faster convergence rate in practice.
Standard ZO-SGD in $\mathbb{R}^{d}$ often incurs a factor of $\sqrt{d}$ in its nonconvex stationarity bound, due to estimator variance. By restricting each layer $\bm{\theta}^{(\ell)}$ to remain near the pre-trained $\bm{\theta}_{0}^{(\ell)}$, DiZO effectively reduces the dimension to $D\!\ll\!d$, improving the rate to $O(\sqrt{D}/\sqrt{T})$. The full proof and detailed analysis are provided in Appendix F.2.
@@ 6 Experiments
Figure 2: Experiments on RoBERTa-large. DiZO outperforms the baselines with and without LoRA. Detailed numbers are presented in Table E.1, and the loss trajectory is shown in Figure E.1.
@@ 6.1 Experimental Settings
Models and datasets. We evaluate DiZO with various models, including medium-sized masked models (liu2019roberta, ) (RoBERTa-large) and large-sized autoregressive models (zhang2022opt, ; touvron2023llama, ) with different size, including OPT-2.7B, OPT-6.7B, OPT-13B, Llama2-7B, Llama3-3B, and Llama3-8B. The total parameter size is ranging from 355M to 13B. We evaluate on the SuperGLUE (wang2019superglue, ) benchmark, MMLU (hendrycks2020measuring, ) and MT-Bench (zheng2023judging, ) benchmarks. More details on datasets are shown in Appendix B.1.
Baseline. We mainly compare with two ZO works, memory-efficient ZO optimization (MeZO) (malladi2023fine, ) and Hessian-informed ZO optimization (HiZOO) (zhao2024second, ). MeZO is a fundamental and representative work in ZO LLM fine-tuning but suffers from slow convergence speed. HiZOO is a recently proposed ZO acceleration for LLM fine-tuning, which leverages the estimated second-order information to speed up. In addition, we also incorporate the parameter-efficient fine-tuning (PEFT) technique LoRA (hu2021lora, ).
Evaluation. For SuperGLUE, we follow prior work (gao2020making, ; malladi2023fine, ), evaluating few-shot and many-shot settings on RoBERTa-large using $k=16$ and $k=512$ samples per class. For each setting, we randomly sample data for training, validation, and testing. For OPT and LLaMA, we use 1000, 500, and 1000 samples respectively. For MMLU and MT-Bench, we adopt the setup in (pan2024lisa, ; luo2024badam, ), fine-tuning on the Alpaca GPT-4 dataset (taori2023stanford, ). All experiments are conducted on NVIDIA A100 and A6000 GPUs, with results averaged over three trials. Full results are provided in Appendix E.
@@ 6.2 Medium-sized Masked Language Models
We conduct experiments on RoBERTa-large across three types of datasets and compare DiZO with two ZO baselines. We also explore PEFT by integrating LoRA. Figure 2 presents the results, while Figure E.1 shows the trajectory of training loss curves, indicating the convergence speed of DiZO and MeZO. Our key findings are as follows:
DiZO greatly increases the convergence speed over MeZO. By using divergence-driven layer adaptation, the loss curve of DiZO decreases much faster, cutting the required iterations by over 50% on SST-2, MNLI, and RTE. Moreover, DiZO improves accuracy by 1.7%, 3.6%, and 8.5%.
DiZO outperforms MeZO and achieves results on par with full fine-tuning. From Figure 2, DiZO consistently surpasses MeZO on all six datasets. Notably, on SST-2 and RTE datasets, DiZO even shows better performance than FO full-parameter fine-tuning, increasing by 0.3% and 1.5%.
@@ 6.3 Large Autoregressive Language Models
Table 3: Experiments results of fine-tuning OPT-2.7B (with 1000 training samples). Better results between MeZO, HiZOO, and DiZO are highlighted in bold.
@@
Table 4: Experiment results on OPT-6.7B (with 1000 training samples).
Dataset
Task Type
SST-2 RTE CB WSC SQuAD
@@
Figure 3: Experiment result on Llama3-3B and Llama3-8B. More results and detailed numbers are shown in Appendix E.4.
To assess the generalizability of DiZO, we run experiments on the OPT and Llama. The overall results are summarized in Table 3, Table 3, and Figure 3 for OPT-2.7B, OPT-6.7B, and Llama series, respectively. We also compare the convergence speeds of DiZO and MeZO on OPT-2.7B across datasets in Figure 4. We highlight key observations from experiments as follows.
DiZO significantly reduces training GPU hours over MeZO. As shown in Table 4, DiZO achieves faster convergence with up to 48% less GPU time by quickly establishing effective layer-wise divergence. Unlike HiZOO, which reduces iterations but suffers from slow throughput due to costly Hessian estimates, DiZO maintains MeZO-level efficiency with a lightweight projection update using only two forward passes.
DiZO outperforms baselines in both standard and parameter-efficient settings. As shown in Table 3, DiZO consistently outperforms MeZO and HiZOO, with or without LoRA, achieving performance close to FO methods. It ranks first on five of seven classification tasks and leads both text generation tasks. These gains extend to OPT-6.7B (Table 3) and Llama models (Figure 3), highlighting the benefit of layer-wise adaptive updates.
Figure 4: Comparison between MeZO and DiZO on convergence iteration, forward pass, and training GPU hours across multiple datasets. Results are presented as proportions, with the percentage of saved GPU hours highlighted for each dataset.
@@ 6.4 Memory and Speed Analysis
Table 5: Memory utilization and speed test on OPT-2.7B on RTE dataset (180 tokens per example on average). : partial gradient-free; DiZO: learning projection with Adam. For a fair comparison, the speed and memory are measured on the same machine with the same settings.
@@
In this section, we examine the memory utilization and convergence speed of DiZO in comparison with both ZO baselines and FO fine-tuning approaches (with and without LoRA). Table 5 presents the results of fine-tuning OPT-2.7B on the RTE dataset, more results are shown in Appendix E.2.
From a memory perspective, DiZO avoids backpropagation and memory-heavy activations, cutting memory use by  90% compared to FO. Its overhead stems only from storing Query and Value weights (16.7% of total). In contrast, HiZOO stores full-layer Hessians, scaling poorly with model size. In terms of convergence speed, DiZO significantly reduces iteration count while maintaining throughput comparable to MeZO, leading to much lower training GPU hours. By comparison, HiZOO achieves less iteration reduction and slows throughput of MeZO by about 1.5× due to Hessian estimation, resulting in only modest savings, or even higher training cost in some cases, such as HiZOO+LoRA on RTE
@@
A notable byproduct of our method is using FO (e.g., with the Adam optimizer) to learn the projections. While this version has memory consumption comparable to LoRA and requires additional training GPU hours, it offers distinct advantages. Since DiZO does not update projections at every iteration, FO-based DiZO exhibits significantly lower average memory usage than FO-based LoRA, with an average memory overhead close to that of the ZO-based DiZO. Although average memory usage may seem less critical in single-process, single-GPU setup, many real-world on-device training scenarios involve multi-process environments (li2024flexnn, ; ye2024asteroid, ). In such cases, the FO-based DiZO can stagger memory usage phases across processes, enabling parallel operations that purely FO methods cannot achieve. Furthermore, compared with ZO-based DiZO, the FO version reduces extra training GPU hours and delivers better performance. These qualities make it particularly appealing for specific on-device training cases.
6.5 Discussion on Potential Alternatives and Limitations
Adaptive learning rate methods may appear analogous to DiZO at first glance, as both introduce per-layer adaptive control over parameter updates, but their principles differ. Methods like Adam and RMSProp adjust update magnitude based on gradient history, controlling how fast parameters move. DiZO, by contrast, uses geometric constraints to guide parameters toward a learnable target distance from a fixed anchor (the pre-trained model), determining where the parameters move. This projection-based approach enables principled, divergence-aware updates that step-size modulation alone cannot replicate. Additionally, adaptive methods maintain gradient moment estimates, adding memory and computational overhead, particularly in LLMs. We empirically compare DiZO to these methods in Appendix D.
Line search could potentially be a simpler method to replace ZO for optimizing projection scalars. However, line search approaches, e.g., backtracking, generally require tuning each layer’s scalar independently, leading to inefficient and unscalable coordinate-wise search. These methods also rely on directional derivatives and assume smooth interactions, which do not generalize well to joint tuning across layers. DiZO avoids these issues by using a ZO-based strategy that updates all scalars simultaneously, stabilized by projection clipping and re-initialization. As shown in Appendix D, replacing ZO with line search under the same forward-pass budget leads to worse performance, confirming its inefficiency in this context.
Limitations. While DiZO demonstrates notable improvements in both accuracy and training efficiency, the theoretical foundations behind its design choices remain incomplete. For example, the choice of using a pre-trained model as the anchor point and the selection of which layer to be projected is primarily supported by empirical observations rather than formal justification(detailed ablation study in Appendix C). The absence of a solid theoretical framework to explain why such design yields consistent performance gains leaves open questions about the optimality of the approach. Nevertheless, our findings offer valuable insights and point to promising directions for future research, particularly in developing anchor-guided, adaptive ZO optimization frameworks with stronger theoretical grounding.